6.28×1013+8.30×1011.

6.28×1013+8.30×1011.
Answer #1

6.28×1013+8.30×1011 = 628x1011+8.30×1011

or, 6.28×1013+8.30×1011 = (628+8.3)x1011

or, 6.28×1013+8.30×1011 = 636.3x1011

or, 6.28×1013+8.30×1011 = 6.363x1013

***********************************************************************************************
This concludes the answers. If there is any mistake or omission, let me know immediately and I will fix it....